### v5

# Efficient Data Selection for Training Genomic Perturbation Models

Abstract
Genomic studies, including CRISPR-based Perturb-seq analyses, face a vast hypothesis space, while gene perturbations remain costly and time-consuming. Gene perturbation models based on graph neural networks are trained to predict the outcomes of gene perturbations to facilitate such experiments. Due to the cost of genomic experiments, active learning is often employed to train these models, alternating between wet-lab experiments and model updates. However, the operational constraints of the wet-lab and the iterative nature of active learning significantly increase the total training time. Furthermore, the inherent sensitivity to model initialization can lead to markedly different sets of gene perturbations across runs, which undermines the reproducibility, interpretability, and reusability of the method. To this end, we propose a graph-based data filtering method that, unlike active learning, selects the gene perturbations in one shot and in a model-free manner. The method optimizes a criterion that maximizes the supervision signal from the graph neural network to enhance generalization. The criterion is defined over the input graph and is optimized with submodular maximization. We compare it empirically to active learning, and the results demonstrate that despite yielding months of acceleration, it also improves the stability of the selected perturbation experiments while achieving comparable test error.
Introduction
Genomic research enables the study of genetic factors underlying various health conditions, opening new avenues for therapeutic development. Techniques such as CRISPR interference (CRISPRi) and PerturbSeq barrangou2016applications have revolutionized the landscape of genomic experimentation by enabling high-throughput screenings. However, the majority of CRISPR-based PerturbSeq experiments are restricted to hundreds of single gene perturbations dixit2016perturb ; adamson2016multiplexed ; peidli2022scperturb due to budget and time constraints, despite having over 20,000 potential gene targets. To assist exploring this vast space, machine learning models have been developed to predict the outcome of gene perturbations on a given cell roohani2024predicting ; bereket2024modelling ; bertin2023recover ; gaudelet2024season ; lopez2023learning ; lotfollahi2023predicting ; geiger2023systematically .
Recent state-of-the-art methods often leverage graph-based models that use gene-gene interaction networks for perturbation prediction roohani2024predicting . These require a substantial number of training samples that stem from genetic experiments pertaining to a specific hypothesis, rendering training costly and time-consuming because it requires feedback from a wet-lab. Active learning has been proposed to address this problem huang2024sequential by defining an iterative interaction between the wet lab and the base model to run experiments for genes that will optimize the training, as shown in Fig. 1(a). This ensures that resources are not wasted in non-informative genetic experiments. However, in this setting, the training can take months, because the process typically has a few iterations ($\leq 5$huang2024sequential and each iteration translates to a CRISPRi-based PerturbSeq experiment that may take 3-5 weeks gasperini2019genome , without accounting for operational delays due to communication or computational hurdles like model retraining.
Another problem with active learning is that it relies on the model to decide the next batch of genes to train on settles2009active . However, the model is initialized randomly, and the first gene selection is by default random, leading to significantly different training sets whenever the model is retrained. This instability introduces several practical challenges. First, the resulting gene sets are less reusable: if different runs select different genes, the collected data may be too model-specific to support downstream analyses and integration with other studies. Second, thrustworthiness is undermined—if selected perturbations vary with initialization, it becomes unclear whether they reflect a true biological signal or model noise. Third, reproducibility suffers, as the same training strategy yields divergent selections in independent experiments. To mitigate this risk, we need methodologies that disentangle the model with the selection strategy. Although density-based active learning is model-independent hacohen2022active , it is not applicable in our context because the genes are not represented numerically (e.g., with features, etc.). Their embeddings are learned through the model roohani2024predicting and active learning acts upon that space huang2024sequential , which renders these selections also model-based.
In the face of these challenges, in this work, we develop a training strategy for gene expression models that is i) significantly faster, ii) operationally simpler, iii) more stable. To this end, we explore data filtering methods kirsch2025active , which select samples before training the model. Such an approach has several advantages that support real-world adaptation:
  • Decrease experimental time. Data filtering selects the full batch of perturbations upfront instead of iteratively. Since Perturb-seq assays are inherently parallelizable, this approach can complete in the time required for a single active learning cycle—resulting in months of real-world acceleration.
  • Reduced operational complexity. Active learning requires repeated coordination between the machine learning and wet-lab teams, model retraining, and acquisition computations. These steps introduce computational, logistic, and communication delays as well as potential errors, which compound across cycles and are diminished with a one-shot data filtering strategy.
  • Independence of model initialization. Since data filtering does not rely on the model, it produces a consistent set of selected genes across different runs. This stability improves reproducibility, trustworthiness, and reusability of the resulting data in downstream analyses.
(a) Active learning.
(b) Data filtering.
Figure 1: The difference between active learning and data filtering for training GEARS roohani2024predicting .
The only search space we can use to define data filtering is the knowledge graph, which is defined between genes based on prior knowledge and is used for message passing by the model. To this end, we propose two graph-based methods to build the train set for one of the state-of-the-art models, GEARS roohani2024predicting , although our methods can be combined with any graph-based predictor. We rely on a theorem for GNN generalization to define a proxy criterion to guide the sample selection based on the input knowledge graph, as seen in Fig. 1(b). The criterion is proven submodular, and it is optimized greedily to build the training set. Our contributions are as follows:
  • Two novel data-filtering criteria (GraphReach and MaxSpec), that select genes from the knowledge graph such that GEARS generalizes better compared to random selection. GraphReach is optimized using a submodular maximization algorithm, and MaxSpec using a greedy heuristic.
  • An empirical comparison including two active learning methods (one of which is state of the art in our problem), indicating that GraphReach achieves at least 5× faster training and increased stability without essentially sacrificing accuracy.
Related Work
The body of related work can be broadly divided into two categories. The first concerns optimal experimental design, where the goal is to select gene perturbations that maximize a downstream causal effect (e.g., a phenotype). The objective is to choose a subset of perturbations (treatments) from a large action space in order to maximize an expected outcome variable, such as T-cell activation  mehrjou2021genedisco . Methods such as Bayesian optimization and online learning have been applied in this context pacchiano2022neural ; lyle2023discobax ; pacchiano2022neural , and are evaluated based on the number of high-reward interventions discovered.
The second branch of relative literature focuses on efficient training strategies for models that predict the gene expression profile following perturbation in single cells. In this setting, the aim is not to optimize a causal phenotype but to train a model that generalizes well to unseen perturbations. This is particularly useful in Perturb-seq experiments, which measure the transcriptomic effects of perturbations via single-cell RNA sequencing. Due to the high cost and long duration of each experimental cycle, active learning methods have been proposed to efficiently select the most informative perturbations for training huang2024sequential .
The closest method to our approach is IterPert huang2024sequential , which uses active learning and prior multimodal knowledge to build a train set for GEARS. Since each iteration can take 3-5 weeks in the wet-lab, the number of iterations is diminished to 5 in contrast to  50 in optimal experimental design mehrjou2021genedisco . The problem is addressed from the perspective of active learning under budget hacohen2022active with the inclusion of prior imaging and perturb-seq studies. In fact, prior multimodal data was so effective, that it produced state of the art results without active learning, i.e., the model IterPert-prior-only. However, such data is not commonly available.
Our methods differ from prior work in several important ways. First, we propose graph-based data filtering approaches that select all perturbations prior to model training, thereby requiring only a single experimental round. Since Perturb-seq experiments can be parallelized, this reduces the experimental duration from 5-6 months to approximately one month. Second, unlike IterPert, our approach relies only on an open-access ontology, making it widely applicable across settings where curated multimodal data is unavailable. Third, by removing model-driven selection, the methods achieve stable and reproducible gene selection. This improves the reusability and interpretability of the resulting data. Finally, one of the proposed methods achieves competitive test error compared to the state-of-the-art active learning method that does not use prior multimodal data, TypiClust huang2024sequential .
Methodology
In this section, we first formulate the problem, clarify the theoretical motivation for our approaches and we introduce the two methods for data filtering.
Formulation
We are given a gene knowledge graph $G=(V,E)$, where $|V|=N$ and $|E|=M$ with adjacency matrix $\mathbf{A}$, and each node represents a gene, some of which correspond to candidate gene perturbations. Our goal is to select a train set $S\subset V$, under a budget constraint $|S|=b$ with $b\ll N$, to experimentally perturb in a wet lab setting.
For each selected gene $u\in S$, the perturbation experiment yields paired measurements of pre- and post-perturbation gene expression: $\mathbf{X}\in\mathbb{R}^{c\times g}$ for $c$ single-cell profiles and $g$ expressed genes (not to be confused with the perturbed genes), and corresponding post-perturbation outputs $\mathbf{Y}_{u}\in\mathbb{R}^{c\times g}$. Thus, the perturbation effect is modeled as an additive change $\hat{P}_{u}\in\mathbb{R}^{g}$ to $\mathbf{X}$ such that $\mathbf{\hat{Y}}_{u}=\mathbf{X}+1_{c}\hat{P}_{u}^{\top}$ where $1_{c}\in\mathbb{R}^{c\times 1}$ is a column vector of ones, and the model learns a mapping $M:V\rightarrow\mathbb{R}^{g}$. Note that the perturbation is run on multiple cells with varying gene expressions.
The model parameters $\theta^{S}$, trained only on the selected set $S$, represent a graph neural network (GNN) such as GEARS. The loss function is the mean squared error between the predicted and the observed expression:
$\mathcal{L}(\mathbf{X},\mathbf{Y},u;\theta^{S})=\|\mathbf{Y}-(\mathbf{X}+1_{c}\hat{P}_{u}^{\top})\|^{2},$ (1)
where $\hat{P}_{u}=M(u;\theta^{S})$. The problem we address is to select the most informative subset of genes to perturb, such that the trained model generalizes well to unseen perturbations $U\subseteq V\setminus S$. Formally, we aim to optimize:
$\underset{S\subset V,\ |S|=b}{\arg\min}\ \sum_{u\in U}\mathcal{L}(\mathbf{X},\mathbf{Y}_{u},u;\theta^{S}).$ (2)
Data Filtering to Maximize Supervision
We now provide relevant literature and an argument based on gradient flow from our train set to motivate the guiding principle of our methods, which is to maximize the set of reachable nodes from our chosen train set.
Since we refrain from utilizing the model’s predictions, we turn to methodology that resembles traditional density-based active learning, such as Core-sets sener2017active , which aims to cover as much of the input space as possible. Core-sets rely on the fact that if we assume each training sample to cover a ball with radius $\delta$ around them, then choosing the training set to maximize coverage of the dataset reduces the generalization error. We also know from the literature that labeling a node results in an information gain for every other node in their receptive field, i.e., $k$-hop neighborhood wang2020unifying ; zhang2022information . Finally, we know that the generalization is improved as we reduce the geodesic distance between the training and the test nodes ma2021subgroup . These findings support our intuition that covering a larger part of the graph with training nodes should be beneficial for the model.
In the case of GEARS, let $\mathbf{H}_{\ell}$ represent the gene embeddings from layer $\ell$ of the simplified graph convolution (SGC) used to predict the perturbation $\hat{P}$. Assuming a single SGC layer for clarity (while omitting subsequent layers between $\mathbf{H}$ and the loss $\mathcal{L}$), we have:
$\mathbf{E}$ $=\mathbf{O}\mathbf{W}_{0},$ (3)
$\hat{\mathbf{A}}$ $=\left(\mathbf{D}^{-1/2}(\mathbf{A}+\mathbf{I}d)\mathbf{D}^{-1/2}\right)^{k},$ (4)
$\mathbf{H}$ $=\hat{\mathbf{A}}\mathbf{E}\mathbf{W}_{1},$ (5)
where $\mathbf{D}$ is the diagonal degree matrix, $\mathbf{W}_{0}\in\mathbb{R}^{N\times d}$ is the embedding lookup table, $\hat{\mathbf{A}}$ is the normalized adjacency with self-loops to the power of $k=1$ (which is the default parameter chosen in the GEARS model), $\mathbf{W}_{1}\in\mathbb{R}^{d\times d^{\prime}}$ are the SGC’s weights, and $\mathbf{O}\in\mathbb{R}^{N\times N}$ is a row-wise one-hot encoding of the candidate gene perturbations. We base our methodology on the fact that message-passing GNNs update only the representations in $\mathbf{W}_{0}$ of nodes that are within the receptive field of the supervised nodes. To formalize this, consider the gradient of the loss $\mathcal{L}$ with respect to the identity embeddings $\mathbf{W}_{0}$ as a function of $\frac{\partial\mathcal{L}}{\partial\mathbf{H}}$:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{0}}$ $=\mathbf{O}^{\top}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{E}}\right),$ (6)
$\frac{\partial\mathcal{L}}{\partial\mathbf{E}}$ $=\hat{\mathbf{A}}^{\top}\,\left(\frac{\partial\mathcal{L}}{\partial\mathbf{H}}\,\mathbf{W}_{1}^{\top}\right).$ (7)
Thus, the full gradient becomes:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{0}}=\mathbf{O}^{\top}\,\hat{\mathbf{A}}^{\top}\,\frac{\partial\mathcal{L}}{\partial\mathbf{H}}\,\mathbf{W}_{1}^{\top}.$ (8)
For an individual embedding row $\mathbf{W}_{0}[i]$ and since $\mathbf{O}$ is one-hot, this expands to:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{0}[i]}$ $=\sum_{j=1}^{N}\hat{\mathbf{A}}[j,i]\cdot\left(\frac{\partial\mathcal{L}}{\partial\mathbf{H}[j]}\,\mathbf{W}_{1}^{\top}\right),$ (9)
which shows that the gradient of $\mathbf{W}_{0}[i]$ depends only on the gradients of its neighbors in $\hat{\mathbf{A}}$. However, $\frac{\partial\mathcal{L}}{\partial\mathbf{H}[j]}=0$ if $j\notin S$. Hence the supervision signal is propagated solely to nodes reachable from the train set, meaning that the identity embedding $\mathbf{W}_{0}$ that is learned for each gene, is updated only if it lies within the receptive field of supervised nodes.
This means that if we choose the train set such that we maximize the reachable nodes, we will maximize in tandem the number of $\mathbf{W}_{0}$’s adjusted to the supervision signal. This is beneficial because, since GEARS is inherently semi-supervised, expanding the supervision allows for the representations of test samples or their neighbors to get updated, which leads to better generalization. This clarifies how covering larger parts of the graph can potentially improve the prediction. Therefore, to maximize the reach of the supervision signal, we should aim to maximize the number of nodes reached by the training set $S$.
Figure 2: An overview of the data filtering methodology. Our sole input is the knowledge graph from GO roohani2024predicting and a list of candidate perturbations. The data filtering algorithm, such as GraphReach selects the set of gene perturbations, and they are given to the wet lab for the experimental part. Finally, the single-cell gene expression data is given to GEARS for training and validation.
GraphReach
We define our data filtering criterion as a function that maximizes the number of nodes reached by the node set $S$, i.e, the receptive field. Specifically for a new node $v$, the criterion is defined as the additional reachability that $v$ adds in the set of train perturbations $S$
$\alpha(v,S)=|R_{L}(S\cup\{v\})-R_{L}(S)|,$ (10)
where $R_{L}$ is the set of nodes reachable from set $S$, based on the number of SGC layers $L$ in the definition of GEARS:
$R_{L}(S)=\{v\in V\mid\mathbf{A}^{\ell}_{uv}>0\text{ for }u\in S,1\leq\ell\leq L\}.$ (11)
The subscript $L$ is omitted in subsequent discussion as it remains stable throughout all experiments. The criterion function selects greedily the node $v$ maximizing $\alpha(v,S)$ in every iteration. This is fundamentally suboptimal baykal2021low ; kirsch2019batchbald , but like many active learning methods lyle2023discobax ; pacchiano2022neural ; tigas2022interventions ; sussex2021near , we will prove our criterion is monotonic and submodular to achieve a guarantee that the result will be at least 11/e1-1/e close to the optimal nemhauser1978analysis .
Lemma 1.
The reachability function $R$ as defined in Eq. 11 is monotonic and submodular.
The reader is referred to the Appendix for the whole proof. Let two sets $S_{t}\subseteq S_{t+1}\subseteq V$ and any node $v\notin S_{t+1}$. Since $S_{t}\subseteq S_{t+1}$, every node reachable from $S_{t}$ is also reachable from $S_{t+1}$, meaning $R(S_{t})\subseteq R(S_{t+1})\Rightarrow~|R(S_{t})|\leq|R(S_{t+1})|$, which proves monotonicity. By definition, adding a node earlier ($\alpha(v,S_{t})$) adds at least as many reachable nodes as adding it later ($\alpha(v,S_{t+1})$), since the nodes in the cut of $S_{t}$ and $S_{t+1}$ can belong to $R(v)$ as well: $R(v)\cap(R(S_{t+1})\setminus R(S_{t}))\geq 0$. This means that any node that appears in $R(v)$ and in $(R(S_{t+1})\setminus R(S_{t}))$ was new for $R(S_{t})$ but is not new for $R(S_{t+1})$. Hence, the additional reachability that $v$ can bring to $S_{t+1}$ compared to the one it can bring to $S_{t}$ is diminished. This sketches out the proof that the reachability is submodular.
The greedy approach requires testing every available perturbation for every addition during a cycle, i.e., $\mathcal{O}(NT)$ where $T$ is the total training budget in samples. We employ a cost effective lazy forward approach to accelerate the acquisition step without loss of accuracy leskovec2007cost . The final data filtering algorithm, called GraphReach from graph reachability, is shown in Alg. 1, and an overview of the overall step-by-step data filtering approach can be seen in Fig. 2.
Algorithm 1 GraphReach
1:budget $B$, graph $G$
2:Train set $S=\emptyset$
3:Criterion $\Delta[v]=\alpha(v,\emptyset)$ for all $v\in V$
4:sort($\Delta$)
5:Heap $\mathcal{L}$ with gene perturbations prioritized by $\Delta$
6:while $\mathcal{L}\neq\emptyset$ and $|S|\leq B$ do
7:  $v$ = $\mathcal{L}[0]$
8:  $u$ = $\mathcal{L}[1]$
9:  $\delta=\alpha(v,S)$
10:  if $\delta>\Delta[u]$ then
11:   $S=S\cup\{v\}$
12:   $\mathcal{L}=\mathcal{L}[1:]$
13:  else
14:   $\Delta[v]=\delta$
15:   sort($\Delta$) and rearrange $\mathcal{L}$ accordingly
16:  end if
17:end while
18:$S$
MaxSpec
Another line of research suggests that GNN generalization improves if the largest singular value (i.e., spectral norm) of the message passing operator is diminishing ju2023generalization . We can translate this as choosing a perturbation such that the underlying graph formed by the nodes that are reachable from $S$, has the smallest possible spectral norm of the respective diffusion matrix. The diffusion matrix is $\mathbf{D}^{-1/2}(\mathbf{A}_{S}+\mathbf{I}d)\mathbf{D}^{-1/2}$ with $\mathbf{A}_{S}=A[R(S),R(S)]$ is the subgraph induced by $R(S)$. Hence we define a new criterion function similar to Eq. 10, but instead of reachability, we utilize the spectral norm as:
$M(\mathbf{A}_{S})=\|\mathbf{D}^{-1/2}(\mathbf{A}_{S}+\mathbf{I}d)\mathbf{D}^{-1/2}\|_{2}$ (12)
We still rely on a cost-effective lazy forward method, hence the algorithm MaxSpec (from maximizing spectral norm) is the same Alg. 1, but substituting function $R$ with $M$.
Experiments
As mentioned in the introduction, data filtering is significantly faster than active learning due to high-throughput genomics platforms like Perturb-seq being explicitly designed for parallelized experiments. This translates to a 5-fold acceleration in our context. Besides speed, here we quantify the changes in other dimensions of the problem and visualize the tradeoff. Specifically, we address these questions:
  • Stability: How much do the proposed genomic experiments change throughout different runs?
  • Test Error: How is the test error of the model affected by the training procedure?
  • Tradeoff: Does the overall tradeoff justify data filtering approaches?
To this end, in this section, we first describe the data used for the experiments, including the gene perturbation datasets and the knowledge graph, the benchmarks used for comparison, as well as the experimental design and the evaluation methods. Afterwards, we showcase the performance of the methods and analyze them to derive conclusions about the soundness of the proposed techniques. The code of the experiments, along with details on computing infrastructure is provided in the supplementary material.
Data
We test our methods in two genomic datasets stemming from PerturbSeq experiments, single-cell CRISPR-based screenings on K562 chronic myeloid leukemia cells:
  • Adamson: A study that investigates the mammalian unfolded protein response of K562 cells to CRISPR interference such as repression. After data curation, it contains 81 binary gene perturbations with approximately 800 cells each adamson2016multiplexed .
  • Replogle: A study that systematically maps genotype-phenotype relationships using CRISPR knockouts. After data curation, it includes 1087 binary gene perturbations with approximately 150 cells each replogle2022mapping .
The datasets are diverse in terms of supervision (150 to 800 cells) and number of perturbations (1087 to 82). We chose them in order to juxtapose the method performances in these two different settings and identify which methods perform adequately in both, taking into consideration the scarcity of relevant datasets.
The graph we focus on for both datasets is based on pathway information from GO51 gene2004gene . Each gene is associated with a number of pathway GO51 terms. The Jaccard similarity between the sets of pathways of two genes, i.e., the fraction of shared pathways, is used to calculate the strength of the edge between them. The graph is sparsified by keeping a predefined number of the most important neighbors for each node. The final graph contains over 9,800 nodes and 200,000 edges, exactly as in roohani2024predicting . It should be noted that GEARS utilizes the whole knowledge graph to learn representations and uses the gene perturbation datasets for supervision only in a semi-supervised manner. The data filtering methods follow suit and utilize the whole graph. For example, the reachability of a gene perturbation in Replogle is not defined only on the 1088 candidate perturbations, but on the whole 9,800 nodes.
(a) Adamson.
(b) Replogle.
Figure 3: The total number of unique genes perturbed in active learning experiments with 3 different random seed initializations. The variation in the data filtering methods is induced solely by different $k$-fold splits. The difference with the other methods corresponds to the sensitivity of the corresponding gene selection process to the model initialization and the first gene selection, which are both random.
Benchmarks
The benchmark models are:
  • Random selection from the available gene perturbations. This is the prevalent practice in application because of its speed and simplicity.
  • ACS-FW pinsler2019bayesian is a Bayesian batch active learning model. It selects perturbations such that the new posterior distribution of the model’s parameters approximates the expected posterior when the whole dataset is available. We utilized the version from the BMDALreg library holzmuller2023framework , which was one of the top-performing methods in similar experiments on active learning for GEARS huang2024sequential .
  • TypiClust hacohen2022active is the state-of-the-art active learning algorithm on our problem that does not rely on multimodal priors huang2024sequential . It has outperformed a plethora (8 to be precise) of models.The algorithm clusters the candidate perturbations based on the final graph layer representation from GEARS. Within each cluster, the typicality is quantified as the inverse of the average distance between each sample and an example’s $K$-nearest neighbors, with $K=20$. The most typical sample is selected.
Note that IterPert does not operate without multi-modal data and hence can not be utilized in our setting.
(a) Adamson.
(b) Replogle.
Figure 4: Tradeoff between training time, predictive error (MSE), and stability of gene selection (the dot size is analogous to mean Jaccard similarity across seeds). Each point corresponds to a different perturbation selection strategy. GraphReach achieves low error and high stability in a single experimental cycle (1 month), compared to model-dependent active learning methods (e.g., TypiClust), which require five cycles and exhibit less stable gene selection.
Design
Similar to  huang2024sequential , we start with a random test set that contains roughly 10% of the available perturbations. The test set stays constant throughout the cycles to ensure there is no interaction between the strategies and the evaluation, which could lead to inconsistencies, e.g., adding a test perturbation that is hard to predict in the train set would diminish the error due to easier evaluation, not due to the strategy’s success. We train the model 5 times sequentially, each time adding to the train set 5% of the available perturbations and keeping 1% for validation. At the end of the process, 30% is used for training and validation, and 10% for testing, leaving 60% of the dataset unutilized. This is imperative to perform an effective comparison between the selected train sets. To be specific, if we increase the train set size, the methods converge unavoidably to very similar final gene selections because the available choices are limited. Hence we choose to leave enough space between choices to highlight the differences between the strategies.
We use a 10-fold cross validation to define the test sets, and perform the whole experiment for 3 different random seed initializations (including model initialization and data splits).We evaluate stability by quantifying how much do the chosen gene perturbations diverge between different seed initialization for each method. To this end, we plot the number of unique genes covered by each method throughout 3 different initializations. Moreover, we quantify the consistency of each method by measuring the overlap between the resulting gene sets using the average pairwise Jaccard similarity defined in the appendix. The evaluation metric for performance is MSE and is computed on the top 20 most differentially expressed genes, i.e., the genes that exhibit the biggest expression differences in the experiment, akin to the literature roohani2024predicting . This is a common practice since the vast majority of the genes have zero expression, meaning the results would be skewed and the differences minuscule had we computed them globally. We use the default GEARS parameters and implementation [^1] https://github.com/snap-stanford/GEARS/tree/master.
Results
Stability
Fig. 3 illustrates the number of unique genes selected to train GEARS for 3 different random initializations and the average Jaccard similarity through all cycles is in Tab. 1 (the full tables can be found in the Appendix). The random seed affects the model and the k-folds of the data. The budget represents the number of genes added per cycle and acts as a baseline. The set of genes retrieved by the data filtering methodologies differs from the budget solely due to different k-fold splits. In contrast, active learning methodologies exhibit near-random variability across different runs. This instability implies that replicating the model would require spending multiple times the original ’gene’ budget, undermining reproducibility. Moreover, the resulting gene sets are tightly coupled to model-specific biases, limiting their reusability for downstream analyses or integration with other studies, and indicating that the selection process is driven more by model bias rather than by a signal in the data.
Data Random
ACS
FW
Typi
Clust
Max
Spec
Graph
Reach
Adamson 0.15 0.15 0.15 0.75 0.75
Replogle 0.10 0.13 0.10 0.78 0.78
Table 1: Average Jaccard similarity on gene selections.
Test Error
As can be seen in Tab. 2, the final model perofrmances of GraphReach are close to TypiClust. Considering the overlapping confidence intervals and similar performance in both datasets, we can deduce that GraphReach performs competitively to the state-of-the-art. In contrast, MaxSpec’s performance in Adamson is inferior. The potential cause is that the Adamson gene sets are limited to 25 nodes, meaning that the spectral norm of the subgraph formed from the respective receptive field might not suffice to act as a proxy for generalization, in contrast to Replogle, where MaxSpec’s performance is strong. Similarly, ACS-FW also performs better in Replogle, possibly due to the increased number of training genes, which makes the model more accurate, as indicated by the MSE. Despite their success in Replogle, both ACS-FW and MaxSpec fall short with the lack of supervision in Adamson, rendering them inferior overall to GraphReach and TypiClust. A final observation is that random selection can be quite effective, albeit with great variance, a fact well known in active learning under budget simeoni2021rethinking ; mittal2019parting ; zhu2019addressing and agreeing with the recent literature on the problem at hand huang2024cost .
Tradeoff
The overall tradeoff between efficiency, test error, and stability is visualized in Fig. 4. We observe that, overall, GraphReach strikes the best balance between low MSE, training efficiency, and stability. We thus deem it a valid alternative to random or active-learning-based training for GEARS, especially given its simple deployment and better overall performance compared to random selection.
Method MSE ($10^{3}$) ↓ AVG
Adamson Replogle
Random $184\pm 13.63$ $159\pm 0.80$ 171.5
ACS-FW $185\pm 9.42$ $154\pm 1.15$ 169.5
TypiClust $180\pm 1.80$ $155\pm 1.21$ 167.5
MaxSpec $192\pm 7.11$ $155\pm 1.47$ 173.5
GraphReach $180\pm 1.93$ $156\pm 1.09$ 168
Table 2: MSE, best is bold and second best is underlined.
Conclusion
Genomic experiments cannot exhaust the available intervention options because they are too costly and time-consuming, thus, GNNs are utilized to predict the outcomes of these interventions (i.e., gene perturbations). However, the training set for a gene expression model consists of genomic experiments themselves, calling for more efficient training strategies. In this work, we proposed graph-based data filtering methods that build the train set without relying on the model’s output. This provides significant real-world speedup (from five months to one) as well as more stable selections. We developed two methods, GraphReach and MaxSpec, which build the train set by selecting the perturbations that are expected to improve the GEARS’ generalization. In the experiments, we observed that GraphReach exhibits competitive test error overall, while being significantly faster and more stable compared to the other methods. In the future, we plan to examine hybrid techniques that utilize data filtering methods to jump-start the training up to a point and blend in active learning for the rest of the runs. Moreover, we plan to experiment with multi-gene perturbations and potentially benchmark with graph active learning methods zhang2021rim ; zhang2021grain .
Appendix
Appendix A Proof of Submodularity
Following the definitions of the main paper, let $R$ be the set function of reachability in the graph, and we have subsets of nodes $S_{t}\subseteq S_{t+1}$. We can start from the definition of submodularity:
$R(S_{t}\cup\{u\})\setminus R(S_{t})\supseteq R(S_{t+1}\cup\{u\})\setminus R(S_{t+1})$ (13)
and continue by reformulating the right and left-hand side of Eq. 13 as follows,
$R(S_{t}\cup\{u\})\setminus R(S_{t})$ $=\big{(}R(S_{t})\cup R(\{u\})\big{)}\setminus R(S_{t})$
$=R(\{u\})\setminus R(S_{t}),$
$R(S_{t+1}\cup\{u\})\setminus R(S_{t+1})$ $=\big{(}R(S_{t+1})\cup R(\{u\})\big{)}\setminus R(S_{t+1})$
$=R(\{u\})\setminus R(S_{t+1}).$
Plugging these reformulations back into Eq. 13 produces,
$R(\{u\})\setminus R(S_{t})\supseteq R(\{u\})\setminus R(S_{t+1}).$ (14)
We shall now derive the reformulated definition of submodularity in Eq. 14 to show that reachability is submodular.
We start with the fact that by definition we have $S_{t}\subseteq S_{t+1}$ and graph reachability $R$ is monotone, consequently:
$\quad R(S_{t})\subseteq R(S_{t+1}).$ (15)
Now recall the anti-monotonicity property of set difference (as illustrated below for sets $X,Y$, and $Z$):
$X\subseteq Y\quad\Longrightarrow\quad Z\setminus X\supseteq Z\setminus Y.$ (16)
So if we add a set difference on Eq. 15 with $R(\{u\})$, we get the desired:
$R(\{u\})\setminus R(S_{t})\supseteq R(\{u\})\setminus R(S_{t+1}).$ (17)
Thus,
$R(S_{t}\cup\{u\})\setminus R(S_{t})\supseteq R(S_{t+1}\cup\{u\})\setminus R(S_{t+1}),$ (18)
which proves that reachability is submodular.
Appendix B Stability with Jaccard Similarity
The stability is defined as the average pairwise Jaccard similarity between the gene sets that each method chooses between runs with different random seed.
$\bar{J}=\frac{1}{\binom{n}{2}}\sum_{1\leq i<j\leq n}J(S_{i},S_{j})=\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}\frac{|S_{i}\cap S_{j}|}{|S_{i}\cup S_{j}|}$ (19)
The full results can be seen in Table 3.
Data Cycle 0 1 2 3 4
Adamson
Random 0.13 0.12 0.14 0.16 0.20
ACS-FW 0.12 0.11 0.14 0.18 0.22
TypiClust 0.12 0.12 0.15 0.18 0.22
MaxSpec 0.76 0.74 0.75 0.75 0.76
GraphReach 0.78 0.76 0.74 0.75 0.76
Replogle
Random 0.04 0.07 0.10 0.14 0.17
ACS-FW 0.04 0.08 0.13 0.18 0.24
TypiClust 0.04 0.07 0.10 0.14 0.18
MaxSpec 0.77 0.78 0.79 0.79 0.79
GraphReach 0.78 0.78 0.79 0.79 0.79
Table 3: Jaccard similarity on gene selections per cycle.
Appendix C Computational Time
The computational time is negligible in our setting, because each sample selection takes less than a minute. That said, GraphReach is around 500 times faster than TypiClust, meaning it scales significantly better with the number of available perturbations. This is important for experiments without predefined candidate perturbations, where the search space can increase from hundreds to tens of thousands.
Appendix D Infrastructure
The computing infrastructure used for the experiments reported in this paper includes a 13th Gen Intel(R) Core(TM) i9-13900K CPU with 24 cores, an NVIDIA GeForce RTX 4070 with 24GB and a CUDA Version: 12.2, and a RAM of 32 GB on an Ubuntu 22.04.